OpenStudy (anonymous):

Write an equation of the line that passes through the point (3,0) and has the slope, m = -2. PLEASE HELP!!!

OpenStudy (anonymous):

HELP!!!

OpenStudy (anonymous):

Use point slope formula: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

so would the answer be y=-2x-3?

OpenStudy (anonymous):

Almost, you need to distrubute the -2 to the -3. \[y-0=-2(x-3)\] \[y=-2x-(-2)3\]

OpenStudy (anonymous):

ooh thanks almost forgot ^.^ thanks

OpenStudy (anonymous):

so its y=-2x+6?

OpenStudy (anonymous):

or is it a -6 instead?

OpenStudy (anonymous):

A negative times a negative will give you a positive. So the positive is correct.

OpenStudy (anonymous):

thanks bunches